Need Help with this problem.
4) A right triangle is a triangle with one angle measuring 90°. In a right triangle the sides are related by Pythagorean Theorem, c² = a² + b², where c is the hypotenuse ( the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 6 feet and 8feet.
Answer:
Show work in this space.

Answer by venugopalramana(3286)

4) A right triangle is a triangle with one angle measuring 90°. In a right triangle the sides are related by Pythagorean Theorem, c² = a² + b², where c is the hypotenuse ( the side opposite the 90° angle). Find the hypotenuse when the other 2 sides’ measurements are 6 feet and 8feet.
Answer:
Show work in this space.
USING ABOVE FORMULA
C^2=6^2+8^2=36+64=100
C=10FEET